OpenStudy (anonymous):

Let f(x)=x^2+5x, g(x)=6-2x, h(x)=2/x-3. Find the following value. 4 g (-2x)=

OpenStudy (cathyangs):

...? How are you supposed to give your answer?

OpenStudy (bahrom7893):

g(-2x) = 6-2*(-2x) = 6+4x 4g(-2x) = 4*(6+4x) = 24+16x

OpenStudy (anonymous):

Thanks bahrom :)

OpenStudy (anonymous):

6 g (-t) = How would I do this one? :/

OpenStudy (bahrom7893):

plug in -t instead of x into g(x) and multiply the whole thing by 6

OpenStudy (anonymous):

Thanks again :)

OpenStudy (anonymous):

36+12t?

OpenStudy (bahrom7893):

yes

OpenStudy (anonymous):

Yee-uh. I got it! Thanks to you